OpenStudy (beautifulsin98):

Solve the system of equations algebraically. Show all of your steps Y= x^2+2x Y= 3x+20

jhonyy9 (jhonyy9):

make y=y so result x^2 +2x = 3x +20 x^2 +2x-3x -20 = 0 x^2 -x -20 = 0 so you can factoriz it completly and will get (x-5)(x+4) = 0 x-5=0 x_1 = 5 x+4=0 x_2 = -4 hope you understand it sure